a 80 mm crankshaft main bearing journal rotates at 3000 rev/min while carring a normal load w= 20kN. If the coefficient of friction in the bearing =0.02, calculate the power lost in friction at this engine speed


1
Expert's answer
2021-04-13T07:02:31-0400

To be given in question

length=80mm

n=3000rev/min

w=20kN

"\mu=0.02"

To be asked in question

Power poss in friction "P_{loss}=?"


We know that



Total friction torque

"l=Rcos\alpha"

"\tau =\frac{2}{3}\mu Wl"

"\tau=\frac{2}{3} \times0.02\times20\times10^3 \times80\times10^{-3}
N\times m"

"P_{loss}=21.33 W"

"P_{loss}=\tau\omega"

"w=2\pi n"

"w=2\pi \times50"

"P_{loss}=21.33\times100\pi"

"P_{loss}=6.697\times10^3 W"